Discuss the working of the adjustment mechanism if, Aggregate Demand (AD) is greater than Aggregate Supply (AS).
OR
If in an economy :
Change in initial investment (ΔI) = ₹ 1,000 crores,
Marginal propensity to save (MPS) = 0•2.
Find the value of the following :
(a) Investment multiplier (K)
(b) Change in final income (ΔY)
1
Expert's answer
2020-12-07T07:07:46-0500

a) Investment multiplier = k = 1/MPS = 1/0.2 = 5

b) (ΔY) = k* ΔI = 5* 1000 = 5000